Compilation of Id^{−}: a subset of Id

Computation Structures Group Memo 315 24 July 1990 Revised 1 November 1990

Zena M. Ariola Arvind

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for this work has been provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-89-J-1988 (MIT) and N0039-88-C-0163 (Harvard).